OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

2x + 3y = M Now move 2x to the right, can you?

OpenStudy (anonymous):

Yeah it is just asking me to show my work.

OpenStudy (anonymous):

okay, so what would it look like if you subtracted 2x from both sides?

OpenStudy (anonymous):

@zacefronfan9 then what stop you from showing your work here?

OpenStudy (anonymous):

@Chlorophyll are you good with interests?

OpenStudy (anonymous):

I have no clue how to do it! hah

OpenStudy (anonymous):

@IStutts well, I'm good at trying ;)

OpenStudy (anonymous):

2x + 3y = M -2x -2x _________________ 3y = M - 2x now try to divide both sides by 3

OpenStudy (anonymous):

@Chlorophyll can I steal you for a few after this then?

OpenStudy (anonymous):

@zacefronfan9 then follow our instruction, then!